OpenStudy (anonymous):

Find all solutions to the equation csc x(2cosx + square root 2)=0 A. x=3pi/4+ 2k(pi) and 7pi/4, where k is any positive integer B. x = 3pi/4+2k(pi) and 5pi/4 , where k is any positive integer C. x = 3pi/4+2k(pi) , where k is any positive integer D. x = 5pi/4+2k(pi) , where k is any positive integer

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=csc+x%282cosx+%2B+square+root+2%29%3D0+ see if got any help

OpenStudy (anonymous):

nope:/

OpenStudy (anonymous):

csc x = 0 1/(sin x) = 0 (multiply by sinx) 1 = 0... no solution 2cosx + √2 = 0 2cos x = -√2 cosx = -√(2)/2... take the arccosine x = 135˚ or 225˚ so the solutions are 135˚, 225˚ plus any integer multiple of 360˚

OpenStudy (anonymous):

in The eqn either coecc x = 0 or cozos x = -1/sqr root of 2 you can solve the graph to see that cos x attains those values at 3pi/4 and 5pi/4 So the correct answer is the second

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

codec x does not attain 0 value